What is the specific heat capacity of a 250, g,


250g object that needs 34, point, 125, k, J,


34.125kJ to have its temperature raised by 65, point, 0, degrees, C,


65.0∘

C?



1
Expert's answer
2021-10-05T10:06:38-0400

The specific heat capacity is given as follows:


"c = \dfrac{Q}{m\Delta T}"

where "Q = 34.125\times 10^{3}J" is the amount of heat given to the object, "\Delta T = 65\degree C" is the temperature change, and "m = 250g = 0.25kg" is the mass of the object. Thus, obtain:


"c = \dfrac{34.125\times 10^3 J}{0.25kg\cdot 65\degree C} = 2.1\times 10^3\dfrac{J}{kg\cdot \degree C}"

Answer. "2.1\times 10^3\dfrac{J}{kg\cdot \degree C}".